# Effects of polar structure and moving ejecta on the dynamics of SGRB jets

### v1

@@ Abstract
At least some short gamma-ray bursts (SGRBs) originate from neutron star mergers, systems that release both a relativistic collimated jet and slower, wider ejecta. These jets evolve through a dense, anisotropic medium shaped by the merger process, resulting in interactions that affect their morphology and observable signatures. We investigate the propagation of SGRB jets through funnel-like structures that can be static or expanding with mildly relativistic speed using 2D axisymmetric relativistic hydrodynamic simulations. Our models are based on radial and angular distributions of density and pressure from general-relativistic magnetohydrodynamic simulations of binary neutron star mergers. We explore different values of the funnel opening angle and density contrast, observing that narrow funnels constrain jet expansion, leading to larger collimation. The polar structure of the ejecta mainly affects the jet evolution in the early stages, whereas the effect of expanding ejecta dominates in later stages. Jets propagating through a low-dense polar funnel move initially faster, while the presence of mildly relativistic ejecta maintains the outflow more collimated after the breakout. Our results highlight the importance of the structure and dynamical properties of the ejecta in shaping SGRB jets and their electromagnetic emission.
@@
pubyear: 2025pagerange: Effects of polar structure and moving ejecta on the dynamics of SGRB jetsEffects of polar structure and moving ejecta on the dynamics of SGRB jets
@@ 1 Introduction
Short gamma-ray bursts (SGRBs) are intense, transient events with characteristic gamma-ray isotropic energies in $\sim 10^{50}-10^{52}$ erg (Gehrels et al., 2009). At least some of them originate from the coalescence of two neutron stars (NSs) (Abbott et al., 2017a). After the merger, a relativistic outflow with an energy of approximately $10^{51}$ erg is released with a typical duration of $t\leq 2$ s (Kouveliotou et al., 1993). The jets can achieve Lorentz factors as high $\Gamma\sim 10^{3}$ (Ghirlanda et al., 2018) and isotropic luminosities between $L_{\rm{iso}}\sim 10^{50}-10^{53}$ erg s-1 (Ghirlanda et al., 2009). These jets exhibit opening angles of approximately $\theta_{\rm{j}}\sim 5^{\circ}-25^{\circ}$ (Rouco Escorial et al., 2023), resulting in intrinsic luminosities of $L_{\rm{j}}\sim 10^{47}-10^{51}$ erg s-1.
@@
An important dynamical process in these systems is the interaction of the relativistic jet with the dynamical ejecta (produced and launched during the merger of the NSs) which in turn produces the prompt gamma-ray emission (Abbott et al., 2017a, b; Goldstein et al., 2017; Hallinan et al., 2017; Margutti et al., 2017; Troja et al., 2017; Lazzati et al., 2017, 2018; Mooley et al., 2018; Ghirlanda et al., 2019). According to observations and general relativistic magnetohydrodynamic (GRMHD) simulations, the ejecta has a mass of $\sim(0.01-0.05)M_{\odot}$ (Cowperthwaite et al., 2017; Smartt et al., 2017; Tanaka et al., 2017; Ruiz et al., 2021), densities of order $\sim(10^{8}-10^{14})$ g cm-3, and magnetic field strengths of $\sim(10^{12}-10^{16})$ G (Ciolfi et al., 2017; Combi & Siegel, 2023; Ruiz et al., 2021). In addition, dynamical ejecta have expanding velocities of the order of $(0.1-0.3)\rm{c}$ (Cowperthwaite et al., 2017; Smartt et al., 2017; Kasliwal et al., 2017) and possibly faster velocities which may reach $0.4\rm{c}$ (Hotokezaka et al., 2013; Rosswog et al., 2024). This interaction creates a complex network of shocks that greatly influences the morphology of the jet and the resulting radiation (Gottlieb et al., 2021; Gottlieb & Globus, 2021; Salafia & Ghirlanda, 2022).
SGRB modeling, primarily examined for the propagation of relativistic and collimated outflows through high-density environments, has been studied analytically (e.g., Begelman & Cioffi, 1989; Bromberg et al., 2011; García-García et al., 2024; Hamidani & Ioka, 2021; Hamidani et al., 2024; Lazzati & Begelman, 2005; Lazzati & Perna, 2019) and numerically (García-García et al., 2023; Gottlieb et al., 2018a, b; Gottlieb & Globus, 2021; Hamidani et al., 2020; Lazzati et al., 2017; Urrutia et al., 2021, 2023). Increasing attention has been paid to realistic and anisotropic environments (e.g., Geng et al., 2019; Mattia et al., 2024; Mpisketzis et al., 2024; Nativi et al., 2022, 2023; Pais et al., 2024; Pavan et al., 2021, 2023, 2025). Advances in GRMHD simulations have revealed the complex structure of this medium, characterized by magnetic fields, velocities gradients, and low-density regions along the symmetry axis, often referred to as “funnels”. These properties are anticipated to play a significant role in the evolution of the SGRB jet (e.g. Bamber et al., 2024; Ciolfi et al., 2017; Ciolfi, 2020; Chabanov & Rezzolla, 2023; Dionysopoulou et al., 2015; Hayashi et al., 2022; Kiuchi et al., 2023; Rosswog et al., 2000; Ruiz et al., 2020).
@@
Previous studies have examined the interaction, evolution, and outcome of an SGRB-like jet when evolving through the dynamical ejecta produced during the merger of two NSs. For example, Murguia-Berthier et al. (2017), examined how a wind (without a funnel) shapes the SGRB jet and determined the conditions under which the relativistic jet can successfully break out the dynamical ejecta through a set of two-dimensional (2D) relativistic hydrodynamic (RHD) simulations. Lazzati & Perna (2019) examined analytically the effect that a uniform and expanding wind has on the evolution of a SGRB-like jet (showing how the expansion of the medium influences the energy profile). Geng et al. (2019) examined through 2.5D relativistic magnetohydrodynamic (RMHD) simulations the time delay influence of an expanding medium with a funnel in the evolution of relativistic jets demonstrating its impact on jet structure and angular distribution. Hamidani & Ioka (2021) analyzed the propagation of an SGRB jet through analytical and semi-analytical models (calibrated with a set of 2D RHD simulations), to study its interaction with an expanding wind without a funnel (as well as a long GRB in the Collapsar scenario), showing how the density and velocity profiles of the medium can influence the morphology of the jet. Gottlieb & Nakar (2022) provide information on the evolution of a magnetized jet propagating through a homologous expanding medium, exploring the evolution and overall dynamics analytically and via three-dimensional RMHD simulations. Pais et al. (2023) and Pais et al. (2024) analyze general relativistic MHD simulations and investigate the effects that a structured and expanding medium has on the dynamics of different SGRB-like jets. The jets vary their luminosity and opening angle while the medium has a fixed funnel. Their findings revealed that the interaction of the medium with the jet influences its dynamics and modifies its structure.
@@
This study aims to investigate the impact of different structured and expanding media on the evolution of a relativistic jet. In order to provide a comprehensive understanding, we study the global effects and also analyze the individual contribution that the structure and expansion of the medium produce. This paper is organized as follows. Section 2 describes the setup and physics in our simulations. The result of the propagation of a relativistic jet through a funnel medium is presented in Section 3. Finally, Section 4 presents the summary and conclusions.
@@ 2 Global setup and models
In order to study the evolution of an SGRB-like jet through different structured and static or expanding media, we run a set of 2D RMHD axisymmetric simulations using the Pluto code in spherical coordinates (Mignone et al., 2012).
Figure 1 illustrates the setup used in our simulations. A relativistic and collimated jet, characterized by its luminosity $L_{\rm{j}}$, opening angle $\theta_{\rm{j}}$, and initial Lorentz factor $\Gamma_{\rm{j,0}}$ propagates through a stratified medium defined by its density $\rho_{\rm{m}}$ and pressure $P_{\rm{m}}$. Due to the modeling in 2D, the jet is axisymmetric. The medium may have an angular structure, including a low-density funnel with an opening angle $\theta_{\rm{f}}$, which provides a channel for the expansion of the jet. Depending on the configuration, the medium can either remain static or expand outward with a constant velocity $v_{\rm{w}}$.
The density of the medium ($\rho_{\rm{m}}$) is composed of four components and is shown in Equation 1. From the GRMHD simulations of (Ciolfi et al., 2017) (with equal NS masses and a soft equation of state, model APR4) the medium has an overall radial profile with $r^{-3}$. An exponential cut-off point was included at $r=r_{\rm{cut}}$ (adapted from Lazzati et al. (2017)). A polar-dependent component (funnel) with opening angle ($\theta_{\rm{f}}$) and a density ratio between the funnel and the medium ($\eta=\rho_{\rm{f}}/\rho_{\rm{m}}$) was incorporated (for further details, see Kawaguchi et al. (2020a, b)). Finally, a background density ($\rho_{\rm{b}}$) was added. The density of the medium $\rho_{\rm{m}}$ reads:
$\rho_{m}(r,\theta)=\frac{\rho_{0}\left(\frac{r}{0.03\ r_{cut}}\right)^{-3}e^{\left(-\frac{r}{r_{cut}}\right)^{3}}}{1+\left(\frac{1}{\eta}-1\right)e^{-10\left(\frac{\theta}{\theta_{f}}\right)^{3}}}+\rho_{b},$ (1)
where $r_{\rm{cut}}=5\times 10^{8}$ cm, $\rho_{0}=3.1\times 10^{8}\rm{g}~{}\rm{cm}^{-3}$, and $\rho_{\rm{b}}=3\times 10^{-3}\rm{g}~{}\rm{cm}^{-3}$. The medium in the computational domain used had a total mass of $\sim 1.1\times 10^{-2}M_{\odot}$.
@@
The jet was set with constant luminosity $L_{\rm{j}}=10^{51}$ erg s-1, $\theta_{\rm{j}}=5^{\circ}$, $\Gamma_{\rm{j,0}}=5$ and the Lorentz factor at infinity of $\Gamma_{\rm{j,\infty}}=400$. The medium could have different funnel opening angles and funnel densities and could be static or expanding at constant velocity $v_{\rm{w}}=0.1\rm{c}$. Two funnel opening angles ($\theta_{\rm{f}}=15^{\circ}$ and $\theta_{\rm{f}}=30^{\circ}$) were considered. The funnel density, defined as a fraction of the medium density, took values of $\eta=1/9,1/6,1/3,1$. The models are named according to their specifics. The acronym “nF” signifies the absence of a funnel, while “F” denotes the presence of a funnel. Similarly, nW” represents a static medium, while W indicates an expanding wind. Furthermore, “l” corresponds to a case in which the funnel is less dense (light, $\eta=1/9$), and “h” indicates that the funnel was denser (heavy, $\eta=1/3$), both compared to their standard funnel correspondent case (F_nW and F_W). The narrow funnel case ($\theta_{\rm{f}}=15^{\circ}$) is denoted by “N”. Table 1 summarizes the parameters used in each model.
Table 1: Model parameters.
@@
F_W 30 1/6 0.1
Except for the jet injection region, at $r=r_{\rm{min}}$ and between $\theta_{\rm{min}}=0^{\circ}$ and $\theta=\theta_{\rm{j}}=5^{\circ}$, the inner boundary had a reflective condition. The rmax boundary had a free outflow condition. The axisymmetric and equatorial boundaries ($\theta_{\rm{min}}$ and $\theta_{\rm{max}}$, respectively) had reflective conditions. The simulations were carried out with a Courant number of $\rm{Co}=0.3$, and the total integration time was $t=0.124$ s. The computational domain extended from $r_{\rm{min}}=2\times 10^{7}$ cm to $r_{\rm{max}}=3.62\times 10^{9}$ cm and between $\theta_{\rm{min}}=0$ and $\theta_{\rm{max}}=\pi/2$. The resolution comprised $N_{r}=30,000$ cells with a logarithmically increasing radial spacing and $N_{\theta}=1,000$ uniformly sized angular cells.

### v2

@@ Abstract
At least some short gamma-ray bursts (SGRBs) originate from neutron star mergers, systems that release both a relativistic collimated jet and slower, wider ejecta. These jets evolve through a dense, anisotropic, and expanding medium produced during the merger process, resulting in interactions that affect their morphology and observable signatures. We investigate the propagation of SGRB jets through funnel-like structures that can be static or expanding with mildly relativistic speed using 2D axisymmetric relativistic hydrodynamic simulations. Our initial conditions are inspired from radial and angular distributions of density and pressure from general-relativistic magnetohydrodynamic simulations of binary neutron star mergers. We explore different values of the funnel opening angle and density contrast. We find that the polar structure of the ejecta mainly affects the jet evolution in the early stages, whereas the effect of expanding ejecta dominates in later stages. Jets propagating through a low-density polar funnel move initially faster, while the presence of mildly relativistic ejecta maintains the outflow more collimated after the breakout. Despite differences in the external medium, the energy dissipation within the jet and cocoon remains similar across models, while the shocked ambient material shows distinct signatures that could be observationally distinguishable. Our results highlight the importance of the structure and dynamical properties of the ejecta in shaping SGRB jets.
@@
pubyear: 2025pagerange: Effects of polar structure and moving ejecta on the dynamics of SGRB jets8
@@ 1 Introduction
Short gamma-ray bursts (SGRBs) are intense, transient events with characteristic gamma-ray isotropic energies in $\sim 10^{49}-10^{52}$ erg (Berger, 2014). At least some of them originate from the coalescence of two neutron stars (NSs) (which our study will be based upon, e.g., Goodman, 1986; Paczynski, 1986; Eichler et al., 1989; Narayan et al., 1992; Abbott et al., 2017). Other SGRB progenitor scenarios include the merger of a NS with a black hole (BH) (e.g., Paczynski, 1991; Narayan et al., 1992; Lee & Kluzniak, 1995; Kluźniak & Lee, 1998; Janka et al., 1999). For a review of the characteristics and progenitors of SGRBs, see (Berger, 2014). After the merger, a relativistic outflow with an energy of approximately $10^{51}$ erg is released with a typical duration of $t_{90}\leq 2$ s (Kouveliotou et al., 1993). The jets can achieve Lorentz factors as high $\Gamma\sim 10^{3}$ (Ghirlanda et al., 2018) and isotropic luminosities between $L_{\rm{iso}}\sim 10^{50}-10^{53}$ erg s-1 (Ghirlanda et al., 2009). These jets exhibit opening angles of approximately $\theta_{\rm{j}}\sim 0.5^{\circ}-25^{\circ}$ (Rouco Escorial et al., 2023), resulting in intrinsic luminosities of $L_{\rm{j}}\sim 10^{45}-10^{51}$ erg s-1.
@@
An important dynamical process in these systems is the interaction of the relativistic jet with the dynamical ejecta (produced and launched during the merger of the NSs) (Murguia-Berthier et al., 2021; Urrutia et al., 2021; Salafia & Ghirlanda, 2022). According to observations and general relativistic magnetohydrodynamic (GRMHD) simulations, the ejecta have a mass of $\sim 10^{-3}-10^{-2}M_{\odot}$ (Cowperthwaite et al., 2017; Smartt et al., 2017; Tanaka et al., 2017; Ruiz et al., 2021), densities of order $\sim(10^{8}-10^{14})$ g cm-3, and magnetic field strengths of $\sim(10^{12}-10^{16})$ G (Ciolfi et al., 2017; Ruiz et al., 2021; Combi & Siegel, 2023). In addition, dynamical ejecta have expanding velocities of the order of $(0.1-0.3)\rm{c}$ (Cowperthwaite et al., 2017; Smartt et al., 2017; Kasliwal et al., 2017), with a fraction reaching velocities exceeding $>0.4c$ (Hotokezaka et al., 2013; Rosswog et al., 2024). Throughout this work, we refer to the expanding ejecta component as the wind. This interaction creates a complex network of shocks that greatly influences the morphology of the jet and the resulting radiation (e.g., Gottlieb et al., 2021; Salafia & Ghirlanda, 2022; Mpisketzis et al., 2024).
SGRB modeling has examined the propagation of relativistic and collimated outflows through environments with masses on the order of $\sim 10^{-2}M_{\odot}$ through both analytical (e.g., Lazzati & Perna, 2019; Hamidani & Ioka, 2021; García-García et al., 2024; Hamidani et al., 2024) and numerical studies (e.g., Lazzati et al., 2017; Gottlieb et al., 2018a, b; Hamidani et al., 2020; Gottlieb & Globus, 2021; Urrutia et al., 2021; García-García et al., 2023). Increasing attention has been paid to modeling anisotropic post-merger environments (e.g., Geng et al., 2019; Pavan et al., 2021, 2023, 2025; Nativi et al., 2022, 2023; Mattia et al., 2024; Mpisketzis et al., 2024; Pais et al., 2024). Recent GRMHD simulations have shown that, during the merger of two compact objects, the combined action of gravitational, centrifugal, and pressure gradients generates a non-uniform environment, often featuring low-density polar regions commonly referred to as “funnels” (Rosswog et al., 2000; Pavan et al., 2021). These structures may improve the jet breakout efficiency by reducing the jet’s baryon loading and energy loss during its propagation through the ejecta (Pavan et al., 2021). Neutrino winds may be present (e.g. Aloy et al., 2005; Murguia-Berthier et al., 2017; Nativi et al., 2021) and may also develop an angularly stratified environment with low-density polar regions due to disk outflows (Perego et al., 2014; Sekiguchi et al., 2016; Hayashi et al., 2022). Both funnel configurations (that shaped by gravitational forces or from neutrino-driven winds) may share a similar morphology, but differ in their origin, composition, and dynamical evolution.
@@
Lazzati & Perna (2019) examined analytically the effect that a uniform and expanding wind has on the evolution of a SGRB-like jet (showing how the expansion of the medium influences the energy profile). Geng et al. (2019) examined through 2.5D relativistic magnetohydrodynamic (RMHD) simulations the time delay influence of an expanding medium with a funnel in the evolution of relativistic jets, demonstrating its impact on jet structure and angular distribution. Hamidani & Ioka (2021) analyzed the propagation of an SGRB jet through analytical and semi-analytical models (calibrated with a set of 2D RHD simulations), to study its interaction with an expanding wind without a funnel (as well as a long GRB in the Collapsar scenario), showing how the density and velocity profiles of the medium can influence the morphology of the jet. Gottlieb & Nakar (2022) provides information on the evolution of a magnetized jet propagating through a homologous expanding medium, exploring the evolution and overall dynamics analytically and via three-dimensional RMHD simulations. Pavan et al. (2021) investigated the propagation of a relativistic jet through BNS merger ejecta, where the ejecta structure was pre-computed using a GRMHD simulation. Their analysis revealed that a pre-existing low-density funnel improves jet breakout efficiency. This funnel structure enables the jet to maintain higher energy by reducing baryon loading and interaction losses during its propagation through the ejecta.
Pais et al. (2023, 2024) analyze GRMHD simulations and investigate the effects that a structured and expanding medium has on the dynamics of different SGRB-like jets. The jets vary their luminosity and opening angle while the medium has a fixed funnel. Their findings revealed that the interaction of the medium with the jet influences its dynamics and modifies its structure.
@@
This study aims to investigate the impact of different structured and expanding media on the evolution of a relativistic jet. In order to provide a comprehensive understanding, we study the global effects and also analyze the individual contributions that the structure and expansion of the medium produce. This paper is organized as follows. Section 2 describes the setup and physics in our simulations. The results of the propagation of a relativistic jet through a funnel medium are presented in Section 3. Finally, Section 4 presents the summary and conclusions.
@@ 2 Global setup and models
In order to study the evolution of an SGRB-like jet through different structured and static or expanding media, we perform a set of two-dimensional axisymmetric RHD simulations using the PLUTO code (v4.4; Mignone et al. 2007) in spherical coordinates. The simulations employ a Harten–Lax–van Leer (HLL; Harten et al., 1983) Riemann solver and a third-order Runge–Kutta time integration scheme (for details see (Gottlieb & Shu, 1996) and (Mignone et al., 2007)).We model the gas evolution assuming a relativistic equation of state with an adiabatic index of 4/3.
Figure 1 illustrates the setup used in our simulations. A relativistic and collimated jet, characterized by its luminosity $L_{\rm{j}}$, opening angle $\theta_{\rm{j}}$, and initial Lorentz factor $\Gamma_{\rm{j,0}}$, propagates through a stratified medium defined by its density $\rho_{\rm{m}}$ and pressure $P_{\rm{m}}$. Due to the modeling in 2D, the jet is axisymmetric. The medium may have an angular structure, including a low-density funnel with an opening angle $\theta_{\rm{f}}$, which provides a channel for the expansion of the jet. Depending on the configuration, the medium can either be approximated as static (if the jet velocity significantly exceeds the medium’s expansion speed) or expand outward with a constant velocity $v_{\rm{w}}$ (since the dynamical ejecta is expelled at mildly relativistic velocities shortly before the merger).
The density of the medium ($\rho_{\rm{m}}$) is composed of four components and is shown in Equation 1. Based on the GRMHD simulations of Ciolfi et al. (2017) (with equal NS masses and a soft equation of state, model APR4), and as in García-García et al. (2023), the density profile of the medium was found to follow a $\rho(r)=\rho_{0}(r/r_{0})^{-3}$ radial profile [^1] The profile was obtained by averaging multiple radial profiles sampled at various polar angles from the 2D meridional density plane of the equal-mass APR4 model presented in Ciolfi et al. (2017). The resulting profile was smoothed, and a $\rho\propto r^{-n}$ profile with an index of $n=2.97\approx 3$ was obtained. (with $r_{0}=1.5\times 10^{7}$ cm and $\rho_{0}=3.1\times 10^{8}\rm{g}~{}\rm{cm}^{-3}$).
An exponential cut-off point, adapted from Lazzati et al. (2017)) was included and set to $r_{\rm{cut}}=5\times 10^{8}$ cm in order for the medium to have a total mass of $1.1\times 10^{-2}M_{\odot}$. We follow the approach of Kawaguchi et al. (2020) to introduce the angular stratification of the environment (where the opening angle of the funnel is set by $\theta_{\rm f}$). We add the constant $\eta$ to define the density ratio between the funnel ($0\leq\theta\leq\theta_{\rm f}$) and the equator ($\theta_{\rm f}\leq\theta\leq\pi/2$). For $\eta=1$ there is no density difference between the funnel and the equator, while for $\eta=9$ the funnel is nine times less dense than the equator. Finally, a background density [^2] The background density is not taken into account when calculating the mass of the medium. ($\rho_{\rm{b}}=3\times 10^{-3}\rm{g}~{}\rm{cm}^{-3}$) was added. The density of the medium $\rho_{\rm{m}}$ reads:
$\rho_{m}(r,\theta)=\frac{\rho_{0}\left(\frac{r}{r_{0}}\right)^{-3}e^{\left(-\frac{r}{r_{cut}}\right)^{3}}}{1+\left(\eta-1\right)e^{-10\left(\frac{\theta}{\theta_{f}}\right)^{3}}}+\rho_{b}.$ (1)
The initial density profile for a static medium with a funnel with $\eta=6$ and $\theta_{f}=30^{\circ}$ is shown in Appendix A. The pressure of the medium is taken to be an ideal gas (with an adiabatic index of $\gamma=5/3$) with a temperature of $T=10^{7}$ K. The medium may be static or may be expanding radially with $v=0.1~{}c$.
@@
[…]
$\rho_{\rm{j}}=\frac{L_{\rm{j}}}{\pi c^{2}r_{\rm{ap}}^{2}v_{\rm{j}}\Gamma_{j,0}\Gamma_{\infty}}$ (2)
where $r_{\rm{ap}}$ is the aperture radius ($r_{\rm{ap}}=r_{\rm{min}}tan\theta_{\rm{j}}$) and $v_{\rm{j}}$ is the jet velocity. Meanwhile, the jet pressure ($p_{\rm{j}}$) is:
$p_{\rm{j}}=\frac{\rho_{\rm{j}}c^{2}}{4}\left(\frac{\Gamma_{\infty}}{\Gamma_{\rm{j,0}}}-1\right).$ (3)
The medium is modeled with varying funnel configurations, characterized by different opening angles $\theta_{\rm{f}}$ and density contrasts $\eta$ (see Figure 6 for one of the density initial conditions that we used).
While motionless environments do not fully capture post-merger conditions, they provide a useful reference to isolate the role of the funnel. Thus, we first explored the effects produced by the funnel; then we examined its interaction with an environment expanding with $v_{\rm{w}}=0.1\,c$. We consider two funnel opening angles, $\theta_{\rm{f}}=15^{\circ}$ and $30^{\circ}$, along with density contrast values of $\eta=1,3,6,9$.
The models are named according to their specifics. The acronym “nF” signifies the absence of a funnel, while “F” denotes the presence of a funnel. Similarly, nW” represents a static medium, while W indicates an expanding wind. Furthermore, “l” corresponds to a case in which the funnel is less dense (light, $\eta=9$), and “h” indicates that the funnel was denser (heavy, $\eta=3$), both compared to their standard funnel correspondent case (F_nW and F_W). The narrow funnel case ($\theta_{\rm{f}}=15^{\circ}$) is denoted by “N”. Table 1 summarizes the parameters used in each model.
Table 1: Summary of model parameters. Each model is characterized by the funnel opening angle ($\theta_{\rm f}$), the density contrast between the polar and equatorial regions ($\eta$), and the expanding velocity of the environment ($v_{\rm w}$). Models labeled with “W” include an expanding wind component, while “nW” indicates a static medium.
@@
F_W 30 6 0.1
Except for the jet injection region, at $r=r_{\rm{min}}$ and between $\theta_{\rm{min}}=0^{\circ}$ and $\theta=\theta_{\rm{j}}=5^{\circ}$, the inner boundary had a reflective condition. The rmax boundary had a free outflow condition. The axisymmetric and equatorial boundaries ($\theta_{\rm{min}}$ and $\theta_{\rm{max}}$, respectively) had reflective conditions. The simulations were carried out with a Courant number of $\rm{Co}=0.3$, and the total integration time was limited to $t=0.124$ s due to computational constraints. The computational domain extended from $r_{\rm{min}}=2\times 10^{7}$ cm to $r_{\rm{max}}=3.62\times 10^{9}$ cm and between $\theta_{\rm{min}}=0$ and $\theta_{\rm{max}}=\pi/2$. The resolution comprised $N_{r}=30,000$ cells with a logarithmically increasing radial spacing and $N_{\theta}=1,000$ uniformly sized angular cells. We extend the computational domain and maintain the spatial resolution employed in García-García et al. (2023) where a comprehensive convergence study was conducted.